OpenStudy (anonymous):

can some help with roots of real number. 1.√ (2x)^8

OpenStudy (anonymous):

Simplify\[(\sqrt{2x})^{2}(\sqrt{2x})^{2}(\sqrt{2x})^{2}\]

OpenStudy (phi):

the idea with simplifying square roots is to get a pair of "the same thing" then you can rewrite \[ \sqrt{a \cdot a} \text{ as } a\] (you take away the square root, and write only 1 of the pair. For example, if you had \[ \sqrt{3^4} \] 3^4 means 3*3*3*3 (3 times itself 4 times) \[ \sqrt{3 \cdot 3 \cdot 3 \cdot 3} \] you can see there are two pairs of 3's (3*3) and (3*3) you can rewrite this as \[ \sqrt{3 \cdot 3 \cdot 3 \cdot 3} =3 \cdot 3\] or if we multiplied everything out , using 3^4= 3*3*3*3=81 and 3*3=9, \[ \sqrt{81} =9 \]

OpenStudy (phi):

For your problem, (2x)^8 is (2x) times itself 8 times. Can you write this as pairs of (2x) ?